Consider a sample of year-end prices for Alphabet, Inc. (Google) over a five year period. Google did not pay a dividend over the sample period.

YEAR PRICE
2012 $570.82
2011 $600.80
2010 $531.93
2009 $337.85
2008 $560.22
2007 $511.91

Calculate the average (arithmetic) return.

Calculate the holding period return

Calculate the geometric return
Answer #1

Please refer to below spreadsheet for calculation and answer. Cell reference also provided.

в с D E -Nm 7 Year 2012 2011 2010 2009 2008 2007 Price 570.82 600.8 531.93 337.85 560.22 511.91 Return (r) -4.99% 12.95% 57.4

Cell reference -

в с DE Year Price 2012 2011 2010 2009 2008 2007 570.82 600.8 531.93 337.85 560.22 511.91 Return (r) =C3/C4-1 =C4/C5-1 =C5/C6-

Hope this will help, please do comment if you need any further explanation. Your feedback would be highly appreciated.